# Axi-majoron for almost everything

### v2

@@ Abstract
The details of the minimal cosmological standard model (MCSM) proposed in Ref. [1] are discussed. The model is based on the scale-symmetry and the global Peccei-Quinn(PQ) symmetry with a key assumption that the latter is broken only in the gravity sector in a scale-invariant manner. We show that the model provides a quite simple unified framework for the unknown history of the universe from inflation to the epoch of big-bang nucleosynthesis, simultaneously addressing key puzzles of high energy theory and cosmology: (i) the origin of scales, (ii) primordial inflation, (iii) matter-antimatter asymmetry, (iv) tiny neutrino masses, (v) dark matter, and (vi) the strong CP-problem. Scale symmetry can be exact, and the Planck scale is dynamically generated. The presence of Gauss-Bonnet term may safely retain dangerous non-perturbative symmetry-breaking effects negligible, allowing a large-field trans-Planckian inflation along the PQ-field. Iso-curvature perturbations of axi-majorons are suppressed. A sizable amount of PQ-number asymmetry is generated at the end of inflation, and conserved afterwards. Domain wall problem is absent due to the non-restoration of the symmetry and the non-zero PQ-number asymmetry. Baryogenesis can be realized by either the transfer of the PQ-number asymmetry through the seesaw sector, or by resonant leptogenesis. Dark matter is purely cold axi-majorons from the mis-alignment contribution with the symmetry-breaking scale of $\mathcal{O}(10^{12}){\rm GeV}$. Hot axi-majorons from the decay of the inflaton become a natural source for a sizable amount of dark radiation. Inflationary gravitational waves have information about the masse parameters of the lightest left-handed and right-handed neutrinos, thanks to the presence of an early matter-domination era driven by the long-lived lightest right-handed neutrino species.
@@ 1 Introduction
In Ref. [1], we proposed a minimal scale-invariant $U(1)_{\rm PQ}$-symmetric extension of the SM [21, 22], dubbed as ‘the minimal cosmological standard model(MCSM)’, to address the aforementioned issues altogether other than the notorious Cosmological Constant(CC) problem. In this work, we show the details of the MCSM. It is a DFSZ-type axion model [18, 19] unified with the singlet majoron model, in which Peccei-Quinn(PQ) field not only provides the axion solution, but it is also responsible for the generation of the masses of right-handed neutrinos. Hence, we call the associated pseudo-Goldstone boson as ‘axi-majoron’. In the sense of unifying axion and majoron, our scenario is similar to the earlier works such as Refs. [20] and [23, 24], based on KSVZ [16, 17] and DFSZ [18, 19] axion models, respectively. Also, in regard of providing a unified framework addressing aforementioned puzzles, one can find works such as Refs. [25, 26]. Our scenario is different from those works in various aspects including dynamical generation of the Planck scale, inflationary physics, generation of matter-antimatter asymmetry and physics of axions. In particular, assuming $U(1)_{\rm PQ}$-symmetry to be broken in the gravity sector only [^3] There are some earlier works such as Refs. [27] and [28] where the PQ symmetry-breaking non-minimal gravitational interaction was considered. In Ref. [27], the authors payed attention only to the impact of such an interaction on the axion-isocurvature perturbations which can be suppressed by a temporary axion mass much larger than the expansion rate during inflation thanks to the interaction. In Ref. [28], initial position of the field was set at the valley of the angular potential for simplicity, removing the impact of angular motion. In both works, the PQ-field are located at the angular minimum position during inflation. In our case, the angular motion and the non-zero angular displacement of the PQ-field at the end of inflation play the critical role for the post inflation cosmology., we show that the dynamics of the inflaton along axi-majoron direction can generate a large amount of PQ-number asymmetry which is eventually converted to the baryon number asymmetry through the seesaw sector. It is a kind of Affleck-Dine(AD) mechanism [29] realized by Peccei-Quinn field. Interestingly, such an asymmetry corresponding to a motion of axi-majoron allows not only AD-mechanism but also spontaneous leptogenesis although the latter turns out to be sub-dominant. Even in the case of PQ-number asymmetry not being enough, our scenario allows resonant leptogenesis [30] to work. Also, very intriguingly, in regard to the axion solution to the strong CP-problem, the symmetry-breaking terms in the gravity sector are harmless at least at perturbative level, and may be so even at non-perturbative level thanks to topological curvature-terms such as Gauss-Bonnet [31]. Iso-curvature perturbations are suppressed, and the domain-wall problem is absent due to the angular motion of PQ-field with non-restoration of the symmetry. Also, a sizable amount of dark radiation from hot axi-majorons can be obtained naturally, and may ameliorate the Hubble tension [32] [33]. The standard thermal bath is recovered by decays of the long-lived lightest right-handed neutrino species. Because of the presence of an early matter domination era, primordial gravitational waves from inflation are expected to have a characteristic spectral feature, containing information of the masses of the lightest left-handed and the lightest right-handed neutrinos. In combinations with dark radiation and inflationary observables, it may provide a unique way to probe such physical quantities. Our scenario provides thus a very minimal, unique unified framework for the unknown history of the universe from inflation to the BBN epoch.
@@
This paper is organized as follows. In Section 2, the model is described. In Section 3, inflation with regard to dynamics of the homogeneous mode, evolutions of perturbations, and reheating are discussed. In Section 4, we show how leptogenesis works in our scenario via various mechanisms thanks to the dynamics of PQ-field. In Section 5, dark matter and dark radiation from axi-majorons are briefly discussed. In Section 6, conclusions are drawn.
@@ 2 The Model
We assume that scale-invariance is a fundamental symmetry and preserved in both of the matter and the gravity sectors, and the Planck scale is dynamically generated by spontaneous breaking of scale symmetry. We also assume that the global Peccei-Quinn symmetry, $U(1)_{\rm PQ}$, is preserved but only in the matter sector, since gravity does not seem to respect global symmetries. Based on this symmetry, in order to realize the axion solution to the strong CP-problem and the seesaw mechanism for neutrino masses and mixings in a minimal unified setup, we consider a scale-invariant Type-II two Higgs doublet extension of the SM [34] [^4] KSVZ-type realization such as Ref. [25] is also possible, and heavy extra quarks can contribute to dark matter relic density in this case. with the seesaw sector in the form of the singlet majoron model [35]. The assignment of PQ-charges to fields is shown in Table 1.
@@
There are four scalar fields as the minimal set: a real scalar $\chi$ which is responsible for generating Planck scale via spontaneous breaking of scale symmetry, Type-II two Higgs doublets, and one complex scalar which plays the role of the Peccei-Quinn field, denoted here as $\Phi$.
@@
Our model is minimal in the sense of the number of fields introduced to accommodate various phenomenological/cosmological/theoretical necessities. It might be compared to the model discussed in Ref. [25] for example. The difference is how the axion solution is realized, either DFSZ-type or KSVZ-type. While our scenario introduces only one extra Higgs doublet as a minimal DFSZ axion model, the compared model introduces two extra heavy quarks as a minimal KSVZ axion model. Esthetically, we think our model is much nicer in regard of its symmetrical structure and the very attractive efficiency achieved in such a simple structure with a minimal set of fields.
Instead of considering all these fields explicitly, we deliver our argument with a simplified version, having only one SM Higgs doublet $H$. The presence of the second Higgs doublet which is essential for a realization of DFSZ axion scenario does not affect the main argument. Paying attention to only these fields and their relevant interactions, we characterize our model by the following action555 The largrangian of the gravity sector can have additional symmetry-breaking terms involving $H_{1}^{\dagger}H_{2}$ such as $\tilde{R}\left(H_{1}^{\dagger}H_{2}\pm{\rm h.c.}\right)$. Also, in the matter sector, there can be terms such as $\Phi^{2}\left(H_{1}^{\dagger}H_{2}\right)+{\rm h.c.}$. We ignore these terms in this work since we will be interested in the case in which the couplings of $H_{1}$ and $H_{2}$ to $\Phi$ are negligible and hence those Higgs fields are likely to stay at the symmetric phase before the electroweak phase transition (EWPT). We will discuss else where the case where initially $H_{1}$ and $H_{2}$ have very large field values during inflation. :
$S_{\rm G}$ $=$ $-\frac{1}{2}\int d^{4}x\sqrt{-{\tilde{g}}}{\tilde{R}}\left[\xi_{\chi}\chi^{2}+2\xi_{h}|H|^{2}+2\xi_{\phi}|\Phi|^{2}\right.$ (2.1)
$\left.+\xi_{+}\left(\Phi^{2}+{\rm c.c.}\right)-i\xi_{-}\left(\Phi^{2}-{\rm c.c.}\right)\right]\ ,$
@@
$S_{\rm M}$ $=$ $\int d^{4}x\sqrt{-{\tilde{g}}}\left[{\tilde{g}}^{\mu\nu}\left(\frac{1}{2}\partial_{\mu}\chi\partial_{\nu}\chi+\partial_{\mu}H^{\dagger}\partial_{\nu}H+\partial_{\mu}\Phi^{\dagger}\partial_{\nu}\Phi\right)-{\tilde{V}}\right]\ ,$ (2.2)
where all $\xi$s are assumed to be real positive definite in order for the effective $M_{\rm P}^{2}$ to be positive in the large field region along any direction, and the potential $\tilde{V}$ is given by
${\tilde{V}}$ $=$ ${\tilde{V}}_{S}+{\tilde{V}}_{F}\ ,$ (2.3)
${\tilde{V}}_{S}$ $=$ $\frac{\lambda_{\chi}}{4}\chi^{4}+\lambda_{h}|H|^{4}+\lambda_{\phi}|\Phi|^{4}$ (2.4)
$-\frac{1}{2}\lambda_{\chi h}\chi^{2}|H|^{2}-\frac{1}{2}\lambda_{\chi\phi}\chi^{2}|\Phi|^{2}-\lambda_{h\phi}|H|^{2}|\Phi|^{2}\ ,$
${\tilde{V}}_{F}$ $=$ $\frac{1}{2}y_{N}\Phi^{*}\overline{\nu_{R}^{c}}\nu_{R}+y_{\nu}\overline{\ell}_{L}{\tilde{H}}\nu_{R}+{\rm h.c.}\ .$ (2.5)
where ${\tilde{H}}\equiv i\tau_{2}H^{*}$ with $\tau_{2}$ being the second Pauli matrix [^6] In the full model, ${\tilde{H}}$ should be replaced to $\tilde{H_{2}}$, the conjugate of the up-type Higgs doublet., and the flavor indices were suppressed. All $\lambda$s in Eq. (2.4) are assumed to be positive real free parameters. Note that, thanks to the Yukawa interaction $y_{N}$, the spontaneous breaking of $U(1)_{\rm PQ}$ is responsible for Majorana masses of RHN fields. For this reason, we will call the associated pseudo-Goldstone boson as axi-majoron.
@@
Decomposing $H$ and $\Phi$ as
$H^{T}(x)=(0,h(x)/\sqrt{2}),\ \ \ \Phi(x)=\phi(x)e^{i\theta(x)}/\sqrt{2}=\left(\phi_{r}(x)+i\phi_{i}(x)\right)/\sqrt{2},$ (2.6)
and introducing a new mass scale $M_{\rm P}$ to be fixed later, the action can be expressed as
@@
$S_{\rm G}$ $=$ $-\frac{1}{2}\int d^{4}x\sqrt{-{\tilde{g}}}{\tilde{R}}M_{\rm P}^{2}\left[\frac{\xi_{\chi}\chi^{2}}{M_{\rm P}^{2}}+\frac{\xi_{h}h^{2}}{M_{\rm P}^{2}}+\frac{\xi_{\phi}\phi^{2}}{M_{\rm P}^{2}}+\frac{\xi_{a}\phi^{2}\cos(2\theta)}{M_{\rm P}^{2}}\right]\ ,$ (2.7)

### v3

# Axi-Majoron : One-shot solution to most of the big puzzles of particle cosmology

@@ Abstract
The details of the minimal cosmological standard model (MCSM) proposed in [The minimal cosmological standard model, arXiv:2403.05390.] are discussed. The model is based on the scale-symmetry and the global Peccei-Quinn(PQ) symmetry with a key assumption that the latter is broken only in the gravity sector in a scale-invariant manner. We show that the model provides a quite simple unified framework for the unknown history of the Universe from inflation to the epoch of big-bang nucleosynthesis, simultaneously addressing key puzzles of high energy theory and cosmology: (i) the origin of scales, (ii) primordial inflation, (iii) matter-antimatter asymmetry, (iv) tiny neutrino masses, (v) dark matter, and (vi) the strong CP-problem. Scale symmetry can be exact, and the Planck scale is dynamically generated. The presence of Gauss-Bonnet term may safely retain dangerous nonperturbative symmetry-breaking effects negligible, allowing a large-field trans-Planckian inflation along the PQ-field. Isocurvature perturbations of axi-Majorons are suppressed. A sizable amount of PQ-number asymmetry is generated at the end of inflation, and conserved afterwards. Domain wall problem is absent due to the nonrestoration of the symmetry and the nonzero PQ-number asymmetry. Baryogenesis can be realized by either the transfer of the PQ-number asymmetry through the seesaw sector, or by resonant leptogenesis. Dark matter is purely cold axi-Majorons from the mis-alignment contribution with the symmetry-breaking scale of $\mathcal{O}(10^{12}){\rm GeV}$. Hot axi-Majorons from the decay of the inflaton become a natural source for a sizable amount of dark radiation. Inflationary gravitational waves have information about the mass parameters of the lightest left-handed and right-handed neutrinos, thanks to the presence of an early matter-domination era driven by the long-lived lightest right-handed neutrino species.
@@ 1 Introduction
In Ref. [20], we proposed a minimal scale-invariant $U(1)_{\rm PQ}$-symmetric extension of the SM [21, 22], dubbed as ‘the minimal cosmological standard model(MCSM)’, to address the aforementioned issues altogether other than the notorious cosmological constant (CC) problem. In this work, we show the details of the MCSM. It is a DFSZ-type axion model [17, 18] unified with the singlet Majoron model, in which Peccei-Quinn(PQ) field not only provides the axion solution, but it is also responsible for the generation of the masses of right-handed neutrinos. Hence, we call the associated pseudo-Goldstone boson as ‘axi-Majoron’. In the sense of unifying axion and Majoron, our scenario is similar to the earlier works such as Refs. [19] and [23, 24], based on KSVZ [15, 16] and DFSZ [17, 18] axion models, respectively. Also, in regard of providing a unified framework addressing aforementioned puzzles, one can find works such as Refs. [25, 26]. Our scenario is different from those works in various aspects including dynamical generation of the Planck scale, inflationary physics, generation of matter-antimatter asymmetry and physics of axions. In particular, assuming $U(1)_{\rm PQ}$-symmetry to be broken in the gravity sector only [^3] There are some earlier works such as Refs. [27] and [28] where the PQ symmetry-breaking nonminimal gravitational interaction was considered. In Ref. [27], the authors payed attention only to the impact of such an interaction on the axion-isocurvature perturbations which can be suppressed by a temporary axion mass much larger than the expansion rate during inflation thanks to the interaction. In Ref. [28], initial position of the field was set at the valley of the angular potential for simplicity, removing the impact of angular motion. In both works, the PQ-field are located at the angular minimum position during inflation. In our case, the angular motion and the nonzero angular displacement of the PQ-field at the end of inflation play the critical role for the post inflation cosmology., we show that the dynamics of the inflaton along axi-Majoron direction can generate a large amount of PQ-number asymmetry which is eventually converted to the baryon number asymmetry through the seesaw sector. It is a kind of Affleck-Dine(AD) mechanism [29] realized by Peccei-Quinn field. Interestingly, such an asymmetry corresponding to a motion of axi-Majoron allows not only AD-mechanism but also spontaneous leptogenesis although the latter turns out to be subdominant. Even in the case of PQ-number asymmetry not being enough, our scenario allows resonant leptogenesis [30] to work. Also, very intriguingly, in regard to the axion solution to the strong CP-problem, the symmetry-breaking terms in the gravity sector in our scenario are harmless at least at perturbative level because of the field-dependent nature, and may be so even at nonperturbative level thanks to topological curvature-terms such as Gauss-Bonnet [31]. Isocurvature perturbations are suppressed, and the domain-wall problem is absent due to the angular motion of PQ-field with nonrestoration of the symmetry. Also, a sizable amount of dark radiation from hot axi-Majorons can be obtained naturally, and may ameliorate the Hubble tension [32] [33]. The standard thermal bath is recovered by decays of the long-lived lightest right-handed neutrino species. Because of the presence of an early matter domination era, primordial gravitational waves from inflation are expected to have a characteristic spectral feature, containing information of the masses of the lightest left-handed and the lightest right-handed neutrinos. In combinations with dark radiation and inflationary observables, it may provide a unique way to probe such physical quantities. Our scenario provides thus a very minimal, unique unified framework for the unknown history of the Universe from inflation to the BBN epoch.
@@
This paper is organized as follows. In Sec. 2, the model is described. In Sec. 3, inflation with regard to dynamics of the homogeneous mode, evolutions of perturbations, and reheating are discussed. In Sec. 4, we show how leptogenesis works in our scenario via various mechanisms thanks to the dynamics of PQ-field. In Sec. 5, dark matter and dark radiation from axi-Majorons are briefly discussed. In Sec. 6, conclusions are drawn.
@@ 2 The Model
We assume that scale-invariance is a fundamental symmetry and preserved in both of the matter and the gravity sectors, and the Planck scale is dynamically generated by spontaneous breaking of scale symmetry. We also assume that the global Peccei-Quinn symmetry, $U(1)_{\rm PQ}$, is preserved but only in the matter sector, since gravity does not seem to respect global symmetries. Based on this symmetry, in order to realize the axion solution to the strong CP-problem and the seesaw mechanism for neutrino masses and mixings in a minimal unified setup, we consider a scale-invariant type II two Higgs doublet extension of the SM [34] [^4] KSVZ-type realization such as Ref. [25] is also possible, and heavy extra quarks can contribute to dark matter relic density in this case. with the seesaw sector in the form of the singlet Majoron model [35].
@@
There are four scalar fields as the minimal set: a real scalar $\chi$ which is responsible for generating Planck scale via spontaneous breaking of scale symmetry, type II two Higgs doublets, and one complex scalar which plays the role of the Peccei-Quinn field, denoted here as $\Phi$. The assignment of PQ-charges to fields is shown in Table 1.
@@
Our model is minimal in the sense of the number of fields introduced to accommodate various phenomenological/cosmological/theoretical necessities. It might be compared to the model discussed in Ref. [25] for example. The difference is how the axion solution is realized, either DFSZ-type or KSVZ-type. While our scenario introduces only one extra Higgs doublet as a minimal DFSZ axion model, the compared model introduces two extra heavy quarks as a minimal KSVZ axion model. Aesthetically, we think our model is much nicer in regard of its symmetrical structure and the very attractive efficiency achieved in such a simple structure with a minimal set of fields.
We characterize our model by the following action:
$S_{\rm G}$ $=$ $-\frac{1}{2}\int d^{4}x\sqrt{-{\tilde{g}}}\mathcal{F}{\tilde{R}}\ ,$ (2.1)
@@
$S_{\rm M}$ $=$ $\int d^{4}x\sqrt{-{\tilde{g}}}\left[{\tilde{g}}^{\mu\nu}\left(\frac{1}{2}\partial_{\mu}\chi\partial_{\nu}\chi+\partial_{\mu}\Phi^{\dagger}\partial_{\nu}\Phi+\sum_{i=1}^{2}\partial_{\mu}H_{i}^{\dagger}\partial_{\nu}H_{i}\right)+\mathcal{L}_{\rm ss}-{\tilde{V}}\right]\ ,$ (2.2)
where the SM Lagrangian for fermions and gauge bosons were omitted for simplicity, $\mathcal{F}\equiv\mathcal{F}_{\rm s}+\mathcal{F}_{\rm sb}$ with
$\mathcal{F}_{\rm s}$$\equiv$ $\xi_{\chi}\chi^{2}+2\xi_{\phi}\left|\Phi\right|^{2}+2\xi_{H_{1}}\left|H_{1}\right|^{2}+2\xi_{H_{2}}\left|H_{2}\right|^{2}\ ,$ (2.3)
$\mathcal{F}_{\rm sb}$$\equiv$ $\xi_{+}\left(\Phi^{2}+{\rm c.c.}\right)-i\xi_{-}\left(\Phi^{2}-{\rm c.c.}\right)+\xi_{H_{+}}\left(H_{1}^{\dagger}H_{2}+{\rm c.c.}\right)-i\xi_{H_{-}}\left(H_{1}^{\dagger}H_{2}-{\rm c.c.}\right)\ ,$ (2.4)
and the seesaw sector Lagrangian and the scalar potential are given by
$-\mathcal{L}_{\rm ss}$ $=$ $\frac{1}{2}y_{N}\Phi^{*}\overline{\nu_{R}^{c}}\nu_{R}+y_{\nu}\overline{\ell}_{L}{\tilde{H}_{2}}\nu_{R}+{\rm H.c.}\ ,$ (2.5)
${\tilde{V}}$ $=$ $\frac{\lambda_{\chi}}{4}\chi^{4}+\lambda_{\phi}|\Phi|^{4}+\lambda_{h_{1}}|H_{1}|^{4}+\lambda_{h_{2}}|H_{2}|^{4}$ (2.6)
-- $\frac{1}{2}\lambda_{\chi\phi}\chi^{2}|\Phi|^{2}-\frac{1}{2}\sum_{I=1}^{2}\left(\lambda_{\chi h_{i}}\chi^{2}+2\lambda_{\phi h_{i}}|\Phi|^{2}\right)|H_{i}|^{2}-\frac{1}{2}\left(\lambda_{\phi h}\Phi^{2}H_{1}^{\dagger}H_{2}+{\rm c.c.}\right)$
-- $\lambda_{12}|H_{1}|^{2}|H_{2}|^{2}-\lambda_{12,\times}|H_{1}^{\dagger}H_{2}|^{2}\ .$
In Eq. (2.5), ${\tilde{H}_{2}}\equiv i\tau_{2}H_{2}^{*}$ with $\tau_{2}$ being the second Pauli matrix, and the flavor indices for $\nu_{R}$ and $\ell_{L}$ were omitted. We assume that the nonminimal gravitational couplings in Eqs. (2.3) and (2.4) are positive definite and field-dependent quantities such as [^5] Such an exponential-field-dependence might be hinting a nonperturbative origin for the couplings.
$\xi_{\alpha}(\chi,\varphi_{\alpha})\equiv\xi_{\alpha,0}\times\exp\left(-\frac{c_{\alpha}\chi}{|\varphi_{\alpha}|}\right)\ ,$ (2.7)
where the subscript ‘α’ stands for all different $\xi$s, and $|\varphi_{\alpha}|=\left(\chi,\left|\Phi\right|,\left|H_{1}\right|,\left|H_{2}\right|,\left|H_{1}^{\dagger}H_{2}\right|^{1/2}\right)$ depending on $\xi_{\alpha}$. For the numerical coefficient $c_{\alpha}$, we assume $c_{\chi}=0$ for $\varphi_{\alpha}=\chi$, otherwise
$\sqrt{\frac{\xi_{\chi,0}\left|\varphi_{\alpha,0}\right|}{M_{\rm P}}}\ll c_{\alpha}\ll\sqrt{\frac{\xi_{\chi,0}}{\xi_{\alpha,0}}}\ ,$ (2.8)
where $|\varphi_{\alpha,0}|$ is the low energy vacuum expectation value of the corresponding field. Then, $\xi_{\alpha}(\chi,\varphi_{\alpha})$ is relevant only for $|\varphi_{\alpha}|\gg c_{\alpha}\chi$ with $\xi_{\alpha}(\chi,\varphi_{\alpha})\simeq\xi_{\alpha,0}$, while turned off for $|\varphi_{\alpha}|\ll c_{\alpha}\chi$ at low energy.
@@
In the theory characterized by the action in Eqs. (2.1) and (2.2), Planck scale at low energy is determined dominantly by the vacuum expectation value (VEV) of $\chi$, since $\xi_{\alpha}|\varphi_{\alpha,0}|^{2}\ll M_{\rm P}^{2}$ for $\varphi_{\alpha}\neq\chi$ is expected. It is then given by
$M_{\rm P}\equiv\mathcal{F}_{0}^{1/2}\simeq\xi_{\chi}^{1/2}\chi_{0}\simeq 2.44\times 10^{18}{\rm GeV}\ ,$ (2.9)
where the subscript ‘0’ stands for the low energy VEV of the associated quantity. The determination of Planck scale implies breaking of scale-invariance. Since there are no explicit symmetry-breaking terms, this should be achieved by spontaneous symmetry-breaking. In Ref. [8], it was shown that Planck scale can emerge naturally in scale-invariant theories even without CW-type dimensional transmutation. This is due to the fact that the conserved scale-current has a kernel which eventually settles down to a constant after some general expansion in a very early epoch, at least well before the last about $60$ inflationary $e$-folds. Scale-anomaly can compromise this picture, but it was shown that, once the renormalization scale, which usually appears as a dimensionful parameter, is replaced by a field, scale symmetry becomes exact, and the existence of a constant kernel remains valid [36]. The constant value of the kernel defines an ellipse in the multidimensional scalar-field space 666 Quantum effects on $\xi_{i}$ can change the elliptic configuration [8], but the impact on our discussion is minor. Hence, we ignore the effects in this work., and eventually determines the Planck scale at low energy, once all fields settle down to their true vacuum positions. This whole picture is still valid even in the case nonminimal couplings depend on fields in such a way of Eq. (2.7).
Without loss of generality, decomposing the scalar fields as
$\Phi=\frac{1}{\sqrt{2}}\phi e^{i\theta}=\frac{1}{\sqrt{2}}\left(\phi_{r}(x)+i\phi_{i}(x)\right)$ (2.10)
and $H_{i}=h_{i}e^{i\theta_{i}}/\sqrt{2}\ (i=1,2)$ for the two electrically neutral Higgs fields, one can express the scalar potential ${\tilde{V}}$ as
${\tilde{V}}(\chi,h,\phi)$ $=$ $\frac{1}{4}\chi^{4}\left[\lambda_{\chi}^{\prime}+\lambda_{\phi}^{\prime}\left({\tilde{\phi}}^{2}-\zeta_{\phi\chi}\right)^{2}+\lambda_{h}\left(\zeta_{1}{\tilde{h}}_{1}^{2}+\zeta_{2}{\tilde{h}}_{2}^{2}-\zeta_{h\phi}{\tilde{\phi}}^{2}-\zeta_{h\chi}\right)^{2}\right.$ (2.11)
-- $\left.\lambda_{\phi h}{\tilde{\phi}}^{2}{\tilde{h}}_{1}{\tilde{h}}_{2}\cos\left(2\theta-\theta_{1}+\theta_{2}\right)\right]\ ,$
where ${\tilde{\phi}}\equiv\phi/\chi$, ${\tilde{h}}_{i}\equiv h_{i}/\chi$, and all $\lambda$s and $\zeta$s in Eq. (2.11) are assumed to be positive real free parameters. When the $\xi_{\chi}$-term in Eq. (2.3) dominates the others in $\mathcal{F}$ at low energy, the potential in Einstein frame, $U\equiv{\tilde{V}}/\Omega^{2}$ with $\Omega\equiv\mathcal{F}/M_{\rm P}^{2}$, can be expressed as
@@
$U({\tilde{h}}_{i},{\tilde{\phi}})$ $\simeq$ $\frac{M_{\rm P}^{4}}{4\xi_{\chi}^{2}}\left[\lambda_{\chi}^{\prime}+\lambda_{\phi}^{\prime}\left({\tilde{\phi}}^{2}-\zeta_{\phi\chi}\right)^{2}+\lambda_{h}\left(\zeta_{1}{\tilde{h}}_{1}^{2}+\zeta_{2}{\tilde{h}}_{2}^{2}-\zeta_{h\phi}{\tilde{\phi}}^{2}-\zeta_{h\chi}\right)^{2}\right.$ (2.12)
-- $\left.\lambda_{\phi h}{\tilde{\phi}}^{2}{\tilde{h}}_{1}{\tilde{h}}_{2}\cos\left(2\theta-\theta_{1}+\theta_{2}\right)\right]\ .$
At the minimum of $U$ satisfying $\partial U/\partial{\tilde{h}}_{i}=0$ for $\tilde{h}_{1,2}$, one finds
$\zeta_{1}\tilde{h}_{1}^{2}$ $=$ $\zeta_{2}\tilde{h}_{2}^{2}=\frac{1}{2}\left[\zeta_{h\chi}+\left(\zeta_{h\phi}+\frac{\lambda_{\phi h}}{4\lambda_{h}\sqrt{\zeta_{1}\zeta_{2}}}\right)\tilde{\phi}^{2}\right]\ .$ (2.13)
The VEV of $\tilde{\phi}$ is then found as